Consider a normal distribution with a mean of 14 and variance of 26. What is the Z score for the value of 16?
Answer #1

Solution :

Given that ,

mean = \mu = 14

variance = \sigma 2 = 26

standard deviation =\sigma  = \sqrt{} \sigma2 = \sqrt{} 26 = 5.10

x = 16

Using z-score formula,

z = x - \mu / \sigma   

z = 16 - 14 / 5.10

z = 0.39